INAF queue observing, 2017 Nov 24-25

Observer: F. Cusano, A. Gargiulo
Telescope Operator: S. Allanson
Support Astronomer: B. Rothberg
Instrument: LUCI, MODS, LBC

Summary

At the sunset the sky is clear.
We started with LUCI, but we had immediately a problem in the mask exchange of LUCI1, LUCI1 will be offline for the rest of the night. At 1:42 we switched to MODS.

Night duration (12\xB0 el twilight) 1:40am- 12:40pm UT = 11 h

Observing time: * *10h (89%)

Weather time loss: * *h (%)

technical time loss * *1h 10min*(11%)*

Data Summary

Project_RUN Objects Instrument Filter ExpTime Notes
standard Feige 110 MODS dualgrating, red grating    
Caccianiga GB6J223554 +341737 MODS red grating 1.17h x 2 (bino) completed
Caccianiga GB6J012126 +034646 MODS red grating 0.3h x 2 (bino) completed
Caccianiga GB6J030251 +215510 MODS red grating 0.33h x 2 (bino) 1 of 3
Antoniucci RY Tau LUCI AO FeII, H, K 1.2h completed
Psf Ao star HD283634 LUCI AO Fe II, H2 etc 6min for RY Tau
Vanzella jdz11 LUCI AO Ks 1h  
Vanzella SL cal LUCI Ks 5min  
Annibali R17 CGCG087 LBC g; r 0.66h 2 of 3 obs
Flat   LUCI AO FeII, H2, Ks, K, H, OH1060   N30 camera
MODS calib   MODS dual and red grating    

Details

1.05 slew to standard HD 228448

  • LUCI1 mask exchange does not work, error message on the mask of LUCI1
  • the SA and D. Thompson are trying to solve the problem

01:42 Switch to MODS, problem with LUCI is persisting

02:02 slewing to Feige 110
  • acqBinoMODs mods1r.20171125.0002-3 ; mods2r.20171125.0002-3
  • 02:09 execBinoMODS, dual grating mods1r.20171125.0004-6, mods1b.20171125.0002-4 ; mods2r.20171125.0004-6,mods2b.20171125.0002-4
  • red grating mods1r.20171125.0007-9 ; mods2r.20171125.0007-9
  • seeing 0.6" on guider at 02:29

02:31 slewing to Caccianiga GB6J223554 +341737
  • acqBinoMODS
  • 02:35 problem in pointing, cancelling the preset, the TO is checking the pointing
  • 02:37 again acqBinoMODS mods1r.20171125.0010-13 ; mods2r.20171125.0010-13

02:52 execBinoMODS GB6J223554 +341737.obs
  • seeing on guider 0.6" at 2:53
  • mods1r.20171125.0014-16 ; mods2r.20171125.0014-16
  • seeing 0.6" on guider at 4:00

04:10 slewing to Caccianiga GB6J012126 +034646
  • acqBinoMODS mods1r.20171125.0017-19 ; mods2r.20171125.0017-19

04:22 execBinoMODS GB6J012126 +034646.obs
  • mods1r.20171125.0020-22 ; mods2r.20171125.0020-22
  • seeing 0.7" on guider at 04:22

04:46 slewing to Caccianiga GB6J030251 +215510
  • acqBinoMODS GB6J030251 +215510.acq mods1r.20171125.0023-25 ; mods2r.20171125.0023-25

04:58 execBinoMODS GB6J030251 +215510.obs
  • mods1r.20171125.0026 ; mods2r.20171125.0026
  • seeing 0.6" on guider at 05:00

05:22 the seeing is stable and very good (0.6-0.7"), we decided to performe LUCI-AO observations using LUCI2, we switch to LUCI2

05:40 executing Antoniucci LUCI-AO target RY Tau, ob ry_luci2.xml
  • seeing 0.6 on guider" at 05:43
  • 5:49 problem with the optical gain; problem solved at 5:56
  • 6:05 the loop is open, problem solved at 06:16
  • 6:24 loop opens, 6:25 started again
  • estimated mag by AO system 9.9 mag, 2x2 Binning, freq 990.98 Hz

07:52 psf standard Antoniucci, HD283634
  • binning 2x2,

08:14 slewing to Vanzella program LUCI-AO, ob VanzellaOBLuci2.xml
  • ref AO stars 13.3 mag, binning 3x3
  • 8:15 pointing problem, checking
  • 8:20 started again
  • 8:41 loop opens, pausing the script. Started again at 8:43
  • 8:51 loop opens, pausing the script
  • 8.55 problem with optical gain and NCPA. Test acquisitions with NCPA OFF for the 191.fits. The image quality is comparable to the previous images, so we continue with NCPA OFF to have the loop stable.
  • fwhm on ref star 0.12"

09:08 sending the script again VanzellaOBLuci2.xml

09:43 sending the script again VanzellaOBLuci2.xml
  • fwhm on of a field star star 0.15"

10:13 seeing limited script of Vanzella, obs VanzellaOBLuci2.xml

10:24 switch to MODS, Grazian I

10:35 problem in the reconfiguration of MODS, secondary probelm, D. Miller is working on it, problem persisting we can not use MODS2, 11:22 switching to LBC

11:22 switching to LBC

11:36 slewing to Annibali 17, CGCG087
  • dohybrid
  • lbcrangebal

11:49 exec CGCG087_1.ob
  • dimm seeing 0.75" at 11:50
  • seeing 0.78" on images at 11:58

12:17 exec CGCG087_2.ob
  • seeing 0.8" on images at 12:29

12:47 end of the night

13:03 starting LUCI2 calibrations (lamp on lamp off sequence)
  • N30 Flat, FeII, luci2.20171125.0260-269
  • N30 Flat, H2, luci2.20171125.0270-279
  • N30 Flat, H, luci2.20171125.0280-289
  • N30 Flat K, luci2.20171125.0290-299
  • N30, Flat Ks, luci2.20171125.0300-309
  • N30, OH_1060 luci2.20171125.0313-322

13:04 starting calibration for MODS
  • flat for standard star - red garting: execMODS --mods1(2) slitflats_5as_r(_mods2).cal mods1r.0027-29 ; mods2r.0036-38
  • arc for standard star/science - dual grating: execMODS --mods1(2) grlamps(_mods2).cal mods1r.0030-32, mods1b.005-7 ; mods2r.0039-41, mods2b.0008-10
  • arc for standard star/science - red grating: execMODS --mods1(2) grlamps_r(_mods2).cal mods1r.0033-35 ; mods2r.0042-44
  • flat for science - dual grating: execMODS --mods1(2) slitflats_1as(_mods2).cal mods1r.0036-41, mods1b.008-13 ; mods2r.0045-50, mods2b.0011-16
  • flat for science - red grating: execMODS --mods1(2) slitflats_1as_r(_mods2).cal mods1r.0042-44 ; mods2r.0051-53
Topic revision: r7 - 05 Jul 2018, FeliceCusano
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback